2x+7y=5
x=19-9y
trying to find a ordered pair

Answer by rfer(16322)

2(19-9y)+7y=5
38-18y+7y=5
-11y=-33
y=3
----------
2x+7*3=5
2x+21=5
2x=-16
x=-8
------------
(-8,3)
--------------
proof
-8=19-9*3
-8=-8